lolokrat:

would the ending matrix not be 5 5 1 1 ?

lolokrat:

lolokrat:

because 3(1)+2(0)=5 3(0)+2(1)=5 1(1)+0(0)=1 1(0)+0(1)=1

karissafrazier:

i'm not 100% sure but i think so

lolokrat:

@vocaloid ?

karissafrazier:

look, i'm sorry im not helpful- i'm trying to be helpful towards ppl who need it

lolokrat:

lol no worries dude i just need clarification from someone who knows how to solve this type of problem

lolokrat:

can yall stop memeing i just need to know if my solution is right lol

Vocaloid:

Multiplying by the identity matrix should *not* change any values of the matrix. The original work is correct. 3(1)+2(0)= 3 + 0, not 3+2 3(0)+2(1)= 0 + 2 not 3 + 2 1(1)+0(0)=1 (Same as original problem) Last cell is not 1(0)+0(1)=1. It is 1*0 + 0*1 = 0 + 0 = 0